• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 07:50:59 on July13,2001

    Entry number 65933

    keyword=shift summary

    01:00 Beam back to hall.
    began with carbon/BEO runs.
    02:20 He-3 target for positron data
    The VDC threshold was set to 0. So all runs before 1983/20983
    are junk. Also, prior to run 1983 Nilanga was making software
    modifications. As a result we could not analyze the right arm
    runs.
    04:20 NMR 48.2%
    04:25 Began to cycle magnets for spectrometer change to 1.72 GeV
    04:35 Beam back
    05:05 Magnets stabalized. Began data taking at 1.72 GeV
    05:15 RF cavity Trip. Beam down
    05:20 Pedestal Run.
    05:20 EPR 48.1%
    05:30 MCC sends "tuned" beam to Hall at 5uA
    05:40 CW Beam Back. Taking positron data at 1.72 GeV.
    Frequent beam down.

    Everything we checked looks o.k. to return to production.
    Noticed shift of scintillator ADC peaks (both arms) and
    Cerenkov PMT #2 and 4 (right arm).